OpenStudy (anonymous):

A dilation maps (2, 3) to (4, 6). a) What is the scale factor of the dilation? b) If (-6, 3) is under the same dilation, what would its new coordinate be?

OpenStudy (perl):

A dilation map has the form: \[\large (x,y)\to (a\cdot x, a\cdot y)\]

OpenStudy (anonymous):

how do I do that

OpenStudy (perl):

notice that each coordinate is being multiplied by a constant

OpenStudy (perl):

how would you go from (2,3) to (4,6) , do you see any relationship between x and y

OpenStudy (perl):

\[\large (2,3)\to (a\cdot 2, a\cdot 3)\]

OpenStudy (anonymous):

multiply

OpenStudy (perl):

right, multiply by what?

OpenStudy (anonymous):

2,3and 4,6

OpenStudy (perl):

what do you multiply each coordinate of x and y

OpenStudy (anonymous):

I don't now see I stuck there

OpenStudy (perl):

\[\large (2,3)\to ({\color{Red} 2}\cdot 2, {\color{Red} 2}\cdot 3)\]

OpenStudy (anonymous):

4,6

OpenStudy (perl):

\[\large (2,3)\to ({\color{Red} 2}\cdot 2, {\color{Red} 2}\cdot 3)=(4,6)\]

OpenStudy (perl):

so the scale factor is 2

OpenStudy (anonymous):

now the part to the B

OpenStudy (perl):

\[\large {(2,3)\to ({\color{Red} 2}\cdot 2, {\color{Red} 2}\cdot 3) \\ (-6,3) \to (2\cdot (-6), 2 \cdot 3)}\]

OpenStudy (perl):

\[\large{ (x,y) \to ({\color{blue} 2}\cdot x, {\color{blue} 2}\cdot y) \\~\\ (-6,3) \to (2\cdot (-6), 2 \cdot 3)} \]